OpenStudy (anonymous):

While Neil Armstrong and Buzz Aldrin walked the Moon, the Apollo 11 command module completed 1 orbit every 119 minutes. Write an expression for the time in hours, h, required to complete n orbits. (Hint: Remember that 60 minutes equals one hour.)

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

Setting up a proportion would work well here.

OpenStudy (anonymous):

yeah i know it is but i don't know how

OpenStudy (anonymous):

Well a proportion is just one fraction set equal to another fraction. One fraction is (1 orbit)/(119 minutes), the other fraction is (n orbits)/(h hours). Convert your minutes to hours and solve for h.

OpenStudy (anonymous):

(Though personally, I would solve first and then do the conversion. Leaving the number of hours as a fraction is more precise than using a decimal approximation.)

OpenStudy (anonymous):

welll the options given to me is either A. 119n=h B.7140n=h C. 119n/60=h D. 60n/119=h. I think it's D but im not sure

OpenStudy (anonymous):

To convert from 119min to hours, divide by 60, so h = (119/60)n

OpenStudy (anonymous):

The Apollo 11 module made 30 orbits. Use the expression you wrote in the previous question to find the number of hours it orbited the Moon. Round your answer to the nearest tenth of an hour. (Enter only the number, such as 29.2 without the units) Is this right?